# Item Cluster-aware Prompt Learning for Session-based Recommendation

### v1

@@ Abstract.
Session-based recommendation (SBR) aims to capture dynamic user preferences by analyzing item sequences within individual sessions. However, most existing approaches focus mainly on intra-session item relationships, neglecting the connections between items across different sessions (inter-session relationships), which limits their ability to fully capture complex item interactions. While some methods incorporate inter-session information, they often suffer from high computational costs, leading to longer training times and reduced efficiency. To address these challenges, we propose the CLIP-SBR (Cluster-aware Item Prompt learning for Session-Based Recommendation) framework. CLIP-SBR is composed of two modules: 1) an item relationship mining module that builds a global graph to effectively model both intra- and inter-session relationships, and 2) an item cluster-aware prompt learning module that uses soft prompts to integrate these relationships into SBR models efficiently. We evaluate CLIP-SBR across eight SBR models and three benchmark datasets, consistently demonstrating improved recommendation performance and establishing CLIP-SBR as a robust solution for session-based recommendation tasks.
@@ 1. Introduction
Despite the progress, most existing SBR approaches primarily focus on the sequential relations between items within a single session (intra-session item relationship), often neglecting the valuable information available across multiple sessions (inter-session item relationship). This significant limitation in inter-session modeling has been identified and proven in multiple studies (Ruocco et al., 2017; Qiu et al., 2020; Wang et al., 2023; Li et al., 2023) as a critical issue in SBRs, hindering the accurate learning of user preferences and potentially affecting recommendation performance. To illustrate, consider three sessions for different users: session one {AirPods Pro, AirPods, AirPods Max}, session two {AirPods, Galaxy Buds, Beats Fit Pro}, and session three {Nike Dri-FIT Running Shirt, Adidas Ultraboost, Beats Fit Pro}. When focusing solely on intra-session information, we can observe that the sessions contain Apple earphones, earbuds, and running gears in each session, respectively. While the intra-session perspective provides immediate sequential insights, it struggles to capture more complex item relationships. For instance, the user of session three might also be interested in Airpods Pro instead of Beats Fit Pro, but it’s challenging for the model to identify this solely from intra-session data, as there are no common items between session one and session three. In contrast, inter-session information can discover the deeper connections by linking Beats Fit Pro across session two and three, and AirPods across session one and two, bridging the gap between session one and three. This highlights the importance of inter-session information, as it enhances the understanding of user preferences beyond what intra-session information alone can achieve.
Although several SBRs attempt to incorporate inter-session item relationships and some have shown effectiveness, most still face limitation in computational efficiency. Specifically, GCSAN (Xu et al., 2019) strives to utilize inter-session information but fails to fully model complex relationships between items, which limits its overall recommendation performance. In contrast, models like GCEGNN (Wang et al., 2020) and LESSR (Chen and Wong, 2020) demonstrate effectiveness in capturing inter-session item relationships. However, they are constrained by high computational complexity, leading to prolonged training times. These limitations underscore the need for a robust framework that can fully exploit complex item relationships across sessions while balancing both effectiveness and efficiency in SBRs.
To address the aforementioned challenges, we propose CLuster-aware Item Prompt learning framework for Session-Based Recomm
@@
endation (CLIP-SBR) that captures complex item relationships and efficiently embeds the information into the SBR models. To overcome the limitation in inter-session modeling, the first module, item relationship mining, constructs a global graph from session data. This graph models both intra- and inter-session item relationships in a single structure. Prior studies (Wu et al., 2019; Xu et al., 2019; Yu et al., 2020; Wang et al., 2020; Chen and Wong, 2020) have demonstrated the effectiveness of graph in capturing item relationships. Inspired by the approaches, the global graph represents items as nodes and item transitions as edges, allowing simultaneous modeling of relationships within and across sessions in one structure. Following graph construction, a community detection method (Traag et al., 2019) is used to identify item clusters that share similar user preferences. These clusters offer valuable insights, as items within the same cluster exhibit close relationships and similar user preferences.
@@
To tackle the limitation in computational efficiency, which is the inefficiency of incorporating complex item relationships, we introduce the second module, item cluster-aware prompt learning. they suffer from high computational complexity. This module enhances the learning capabilities of SBR models efficiently by integrating learnable soft prompts for each identified item cluster. Prompt learning (Liu et al., 2023b), originally introduced in natural language processing, began as a way to provide large language models with task-related information through natural language instructions, known as prompts. The concept evolved into learnable prompts, with discrete natural language prompts termed hard prompt and continuous, trainable prompts termed soft prompts. Soft prompts have demonstrated significant success due to their ability to provide task-related cues to pre-trained models during the tuning phase (Li and Liang, 2021; Shin et al., 2020; Lester et al., 2021). This approach has expanded to fields such as computer vision (Jia et al., 2022; Zhou et al., 2022; Ge et al., 2023) and graph representation learning (Sun et al., 2022, 2023; Liu et al., 2023a; Fang et al., 2024), where soft prompts are used to deliver task-specific and data-related information during training. Notably, GPF (Fang et al., 2024) and SUPT (Lee et al., 2024) have proven the ability of soft prompt to deliver data-related information during tuning process. Inspired by the previous studies, we adopt soft prompts to integrate mined item relationship information into SBR models during training. By tailoring learnable soft prompts to specific item cluster, SBR models can effectively and efficiently learn user preferences embedded in session data, capturing both intra- and inter-session relationships. This comprehensive approach not only enhances recommendation performance in an efficient way but also ensures adaptability, making the CLIP-SBR framework universal across various SBR models. We examine CLIP-SBR on eight SBR baselines and three benchmark.
@@
  • We propose a novel approach to enhance SBRs by incorporating item cluster-aware prompt into the model training phase.
@@
  • We conduct extensive experiments on five representative SBRs and three benchmark datasets to demonstrate the superiority of the proposed CLIP-SBR framework.
@@ 2.1. Session-based Recommendation
Session-based Recommendation (SBR) predicts the next user interaction based on recent behavior sequences, effectively capturing short-term user preferences. Recent advancements in deep learning have significantly enhanced SBRs by enabling more sophisticated modeling of item transitions within sessions. GRU4Rec (Hidasi et al., 2016) introduced Gated Recurrent Units (GRUs) for sequential dependency modeling, while models like NARM (Li et al., 2017) and STAMP (Liu et al., 2018) leverage attention mechanisms to capture user intent and preference. CORE (Hou et al., 2022) addresses prediction inconsistencies by unifying session and item embeddings, reducing representation gaps. Graph neural networks (GNNs) have further improved SBRs by representing session sequences as graphs, as seen in SRGNN (Xu et al., 2019), which aggregates item information through gated GNNs, TAGNN (Yu et al., 2020) which employs target-aware attention to incorporate user interests. While these methods effectively model intra-session relationships, they often overlook the valuable inter-session information, limiting their ability to fully understand complex user behaviors. To address this, several inter-session SBRs have been proposed. GCSAN (Xu et al., 2019) uses self-attention on session graphs to capture intricate item relationships, while GCEGNN (Wang et al., 2020) combines item-level and session-level information using graph convolution and self-attention mechanisms. LESSR (Chen and Wong, 2020) introduces an edge-order preserving aggregation scheme and a shortcut graph attention layer to effectively capture global item relationships. Despite their effectiveness, they either inadequately capture inter-session information (GCSAN) or are computationally intensive (GCEGNN, LESSR). Our work proposes incorporating both intra- and inter-session item relationships using efficient cluster-aware prompts to enhance SBR performance.
@@ 2.2. Prompt Learning
Prompt learning is initially introduced in natural language processing (NLP) as a novel method to guide pre-trained language models in performing specific tasks by providing task-related information (Liu et al., 2023b). It originated with the use of discrete hard prompts: text-based instructions designed to provide task related information to pre-trained models (Brown, 2020). However, due to the limitations of hard prompts that requires human intervention and limited adaptability, the concept has evolved to include continuous soft prompts. Soft prompts are learnable vectors that do not directly correspond to natural language but are optimized during training to enhance model performance on a given task (Li and Liang, 2021; Shin et al., 2020; Lester et al., 2021). In recent years, the effectiveness of soft prompts has been demonstrated beyond NLP, extending to domains such as computer vision (CV) and graph representation learning. In the CV field, visual soft prompts have been well-studied and shown great potential in cross-task generalization, domain adaptation, and visual-language models (Ge et al., 2023; Zhou et al., 2022; Jia et al., 2022). Similarly, in the graph field, pioneering works (Sun et al., 2022; Liu et al., 2023a) align task-specific soft prompts and downstream tasks, such as link prediction. GPF (Fang et al., 2024) and SUPT (Lee et al., 2024) further enhance by introducing soft prompts to every node, therby providing graph data-related information in tuning phase, enabling better model adaptation to complex graph structures and tasks. As a result, prompt learning has emerged as a versatile tool that can efficiently improve task performance by providing task- or data-related information across various fields.
@@ 3. Proposed Method
The CLIP-SBR aims to exploit both intra- and inter-session item relationships to effectively model the user preferences embedded in sessions and enhance existing SBRs with the information. Figure 1 presents the overall flow of CLIP-SBR, which comprises two main components: 1) Item Relationship Mining. This component constructs a graph structure from session data to model item relationships with sufficient depth. It then identifies item clusters based on sequential dpendencies by employing existing community detection method. 2) Item Cluster-aware Prompt Learning. This component integrates learnable prompt vectors into item embeddings before they are fed into SBR models. The prompt vectors provide SBR models with information about item clusters during training. In this chapter, we first present the problem statement and then introduce the components of CLIP-SBR in detail.
@@ 3.3. Item Cluster-aware Prompt Learning
Table 2. Comparison of MRR@5, Recall@5 and average improvement (%) on the three datasets. Our proposed models are highlighted in gray. The best results are highlighted in boldface, and the best results among intra-session models are underlined.
Category Model Last.fm Xing Reddit
MRR@5 Recall@5 Imp.(%) MRR@5 Recall@5 Imp.(%) MRR@5 Recall@5 Imp. (%)
GRU4Rec 7.70 13.39 - 10.33 15.71 - 35.45 46.40 -
CLIP-GRU4Rec 10.07 17.24 +29.77 11.12 16.90 +7.61 40.21 50.47 +11.10
NARM 5.50 9.74 - 9.66 15.41 - 33.08 44.48 -
CLIP-NARM 8.10 13.96 +45.30 12.04 17.81 +20.11 38.50 48.79 +13.04
intra- CORE 10.51 17.70 - 10.27 17.79 - 29.33 46.54 -
session CLIP-CORE 10.67 18.11 +1.92 12.65 19.59 +16.65 31.33 47.56 +4.51
SRGNN 9.02 14.66 - 14.03 19.37 - 38.04 47.75 -
CLIP-SRGNN 12.12 18.87 +31.54 14.18 19.72 +1.44 39.70 49.73 +4.26
TAGNN 9.25 15.30 - 12.68 18.37 - 40.07 51.10 -
CLIP-TAGNN 12.82 19.77 +33.91 14.70 21.07 +15.31 41.22 52.01 +2.33
GCSAN 4.47 8.69 - 13.02 16.82 - 36.03 44.13 -
CLIP-GCSAN 7.05 11.59 +45.54 15.65 20.34 +20.56 39.07 48.21 +8.84
inter- GCEGNN 10.81 17.32 - 12.83 19.41 - 38.17 48.38 -
session CLIP-GCEGNN 11.05 17.60 +1.92 14.36 21.58 +11.55 37.93 48.19 -0.51
LESSR 7.06 11.05 - 14.67 18.96 - 39.05 47.90 -
CLIP-LESSR 6.85 10.96 -1.89 14.94 19.07 +1.21 39.07 48.21 +0.35
@@ 4. Experiments
  • RQ3: Does the Item Cluster-aware Prompt Learning module effectively and efficiently integrate item relationships into the recommendation process effectively and efficiently?
@@ 4.1.2. Baselines
To evaluate the effectiveness of our approach, we compare its performance against a diverse set of well-established methods. The first five baselines focus soley on intra-session information, while the last three incorporate both intra-session and inter-session information. We assess our method across various model architectures, including those based on RNNs, GNNs, and attention mechanisms, to demonstrate the applicability and effectiveness of our method across different model architectures.
@@
  • LESSR (Chen and Wong, 2020) introduces shortcut connections between items within a session and incorporates sequence information into graph convolution with the help of GRU.

### v2

@@ Abstract.
Session-based recommendation (SBR) captures dynamic user preferences by focusing on item sequences within sessions. However, many existing methods primarily model intra-session item relationships (within a single session) while overlooking inter-session relationships (between items across different sessions), limiting their effectiveness to fully capture complex item relations. Although several studies incorporate inter-session information, they often face high computational complexity, leading to prolonged training times and reduced efficiency. To address these issues, we propose the CLuster-aware Item Prompt learning framework for Session-Based Recommendation (CLIP-SBR) that consists of two modules: 1) item relationship mining that incorporates a global graph to extract both intra- and inter-session item relationships effectively from session data, and 2) item cluster-aware prompt learning which utilizes soft prompts—lightweight learnable parameters to integrate the mined item relationships into SBR training efficiently. We validate CLIP-SBR through experiments on eight SBR models and three benchmark datasets. Results show consistent improvements in recommendation performance, demonstrating CLIP-SBR’s potential as a robust framework for SBRs.
@@ 1. Introduction
Despite the progress, most existing SBR approaches primarily focus on the sequential relations between items within a single session (intra-session item relationship), often neglecting the valuable information available across multiple sessions (inter-session item relationship). This significant limitation in session modeling has been identified and proven in multiple studies (Ruocco et al., 2017; Qiu et al., 2020; Wang et al., 2023; Li et al., 2023) as a critical issue in SBRs, hindering the accurate learning of user preferences and potentially affecting recommendation performance. Consider three sessions: session one contains AirPods Pro, AirPods, AirPods Max, session two includes AirPods, Galaxy Buds, Beats Fit Pro, and session three focuses on unrelated items like running gear and Beats Fit Pro. Intra-session modeling captures sequential patterns but misses broader connections. Inter-session modeling, however, identifies that Beats Fit Pro in session three links to session two, which in turn connects to session one via AirPods. This enables the model to bridge session one and three, revealing deeper item relationships. For instance, the user of session three might also be interested in Airpods Pro instead of Beats Fit Pro, but it’s challenging for the model to identify this solely from intra-session data, as there are no common items between session one and session three. In contrast, inter-session information can discover the deeper connections by linking Beats Fit Pro across session two and three, and AirPods across session one and two, bridging the gap between session one and three. This highlights the importance of inter-session information, as it enhances the understanding of user preferences beyond what intra-session information alone can achieve.
Although several SBRs attempt to leverage inter-session relationships, they often face efficiency challenges. For example, GCSAN (Xu et al., 2019) partially incorporates inter-session information but struggles with limited modeling capabilities, while GCEGNN (Wang et al., 2020) effectively captures complex relationships but incurs high computational costs. These challenges highlight the need for a framework that balances both effectiveness and efficiency.
@@
To address the challenges, we propose CLuster-aware Item Prompt learning framework for Session-Based Recommendation (CLIP-SBR) that captures complex item relationships and efficiently embeds the information into the SBR training. To overcome the limitation in session modeling, the first module, item relationship mining, constructs a global graph from session data. This graph models both intra- and inter-session item relationships in a single structure. Prior studies (Wu et al., 2019; Xu et al., 2019; Yu et al., 2020; Wang et al., 2020; Chen and Wong, 2020) have demonstrated the effectiveness of graph in capturing item relationships. Inspired by the approaches, the global graph represents items as nodes and item transitions as edges, allowing simultaneous modeling of relationships within and across sessions in a structure. Following graph construction, a community detection method (Traag et al., 2019) is used to identify item clusters that share similar user preferences. The clusters offer valuable insights, as items within a cluster exhibit close relationships and similar user preferences.
@@
To tackle the limitation in efficiency caused by the challenge of incorporating complex item relationship, we introduce the second module, item cluster-aware prompt learning. This module enhances the learning capabilities of SBRs efficiently by integrating learnable soft prompts for each identified item cluster. Prompt learning (Liu et al., 2023b), provides large language models with task-related information through natural language instructions, known as prompts. The concept evolved into learnable prompts, with discrete natural language prompts termed hard prompt and continuous, trainable prompts termed soft prompts. Soft prompts have demonstrated significant success due to their ability to provide task-related cues to pre-trained models during the tuning phase (Li and Liang, 2021; Shin et al., 2020; Lester et al., 2021). This approach has expanded to fields such as computer vision (Jia et al., 2022; Zhou et al., 2022; Ge et al., 2023) and graph representation learning (Sun et al., 2022, 2023; Liu et al., 2023a; Fang et al., 2024), where soft prompts are used to deliver task-specific and/or data-related information during tuning. GPF (Fang et al., 2024) and SUPT (Lee et al., 2024) have proven the ability of soft prompt to deliver data-related information during learning process. Inspired by the previous studies, we adopt soft prompts to integrate mined item relationship information into SBR models during training. By tailoring learnable soft prompts to specific item cluster, SBR models can effectively and efficiently learn user preferences from sessions, capturing both intra- and inter-session relationships.
Unlike previous approaches, our method unifies intra- and inter-session item relationships within a global graph, capturing complex relationships across sessions while preserving local patterns. This enables effective extraction of reusable prior knowledge that enhances SBR training without repeated graph construction or clustering. Furthermore, CLIP-SBR is the first to integrate soft prompts directly into SBR training, leveraging cluster-aware prompts to improve recommendation performance while ensuring adaptability across diverse models, positioning it as a universal framework for session-based recommendation.
@@
  • We propose a novel approach to enhance SBRs by incorporating cluster-aware item prompt into the model training phase.
@@
  • We conduct extensive experiments on eight representative SBRs and three benchmark datasets to demonstrate the superiority of the proposed CLIP-SBR framework.
@@ 2.1. Session-based Recommendation
Session-based Recommendation (SBR) predicts the next user interaction based on recent behavior sequences, effectively capturing short-term user preferences. Recent advancements in deep learning have significantly enhanced SBRs by enabling more sophisticated modeling of item transitions within sessions. GRU4Rec (Hidasi et al., 2016) introduced Gated Recurrent Units (GRUs) for sequential dependency modeling, while models like NARM (Li et al., 2017) and STAMP (Liu et al., 2018) leverage attention mechanisms to capture user intent and preference. CORE (Hou et al., 2022) addresses prediction inconsistencies by unifying session and item embeddings, reducing representation gaps. Graph neural networks (GNNs) have further improved SBRs by representing session sequences as graphs, as seen in SRGNN (Xu et al., 2019), which aggregates item information through gated GNNs, TAGNN (Yu et al., 2020) which employs target-aware attention to incorporate user interests. SCL (Shi et al., 2023) optimizes the item representation with contrastive learning to improve effectiveness of SBRs.
While these methods effectively model intra-session relationships, they often overlook the valuable inter-session information, limiting their ability to fully understand complex user behaviors. To address this, several SBRs have been proposed to better model item relationships. GCSAN (Xu et al., 2019) utilizes self-attention on session graphs to capture global intricate item relationships, while GCEGNN (Wang et al., 2020) combines item-level and session-level information using graph convolution and self-attention mechanisms. HADCG (Su et al., 2023) adopts hyperbolic geometry to extract hierarchical patterns and capture both intra- and inter- session item relationships. Despite their effectiveness, they either struggle to fully capture complex item relationships, which limits their ability to model user preferences and impacts recommendation performance, or suffer from high computational complexity, resulting in prolonged training times. Our work proposes incorporating both intra- and inter-session item relationships using efficient cluster-aware prompts to enhance SBR performance.
@@ 2.2. Prompt Learning
Prompt learning is initially introduced in natural language processing (NLP) as a novel method to guide pre-trained language models in performing specific tasks by providing task-related information (Liu et al., 2023b). It originated with the use of discrete hard prompts: text-based instructions designed to provide task related information to pre-trained models (Brown, 2020). However, due to the limitations of hard prompts that requires human intervention and limited adaptability, the concept has evolved to include continuous soft prompts. Soft prompts are learnable vectors that diverge from direct natural language representations and are instead optimized during training to enhance model performance on a given task (Li and Liang, 2021; Shin et al., 2020; Lester et al., 2021). In recent years, the effectiveness of soft prompts has been demonstrated beyond NLP, extending to domains such as computer vision (CV) and graph representation learning. In the CV field, visual soft prompts have been well-studied and shown great potential in cross-task generalization, domain adaptation, and visual-language models (Ge et al., 2023; Zhou et al., 2022; Jia et al., 2022). Similarly, in the graph field, pioneering works (Sun et al., 2022; Liu et al., 2023a) align task-specific soft prompts and downstream tasks, such as link prediction. GPF (Fang et al., 2024) and SUPT (Lee et al., 2024) further enhance by introducing soft prompts to every node, therby providing graph data-related information in tuning phase, enabling better model adaptation to complex graph structures and tasks. As a result, prompt learning has emerged as a versatile tool that can efficiently improve task performance by providing task- or data-related information across various fields.
@@ 3. Proposed Method
CLIP-SBR leverages both intra- and inter-session item relationships to effectively capture user preferences embedded in sessions, enhancing the training of existing SBR models with this enriched information. Figure 1 presents the overall flow of CLIP-SBR framework, which comprises two main components: 1) Item Relationship Mining which constructs a graph from session data to capture item relationships and identifies item clusters using a community detection method. 2) Item Cluster-aware Prompt Learning that integrates learnable prompt vectors into item embeddings before they are fed into SBR models. The prompt vectors provide SBR models with information about item clusters during training. In this chapter, we first present the problem statement and then introduce the components of CLIP-SBR in detail.
@@ 3.2.1. Graph Construction
Algorithm 1 The whole procedure of CLIP-SBR
Input: Sessions $S$, Items $V$
Output: Recommendation lists
1:  for each session $\mathcal{S}_{t}\in\mathcal{S}$ do
2:     Construct session graph $\mathcal{G}_{t}$ for session $s$;
3:  end for
4:  Aggregate session graphs to form global graph $\mathcal{G}_{g}$;
5:  Get $Partition$ by applying $Leiden$ algorithm on $\mathcal{G}_{g}$ via Eq. (1);
6:  for each iteration do
7:     Get initial item embeddings $\mathbf{V}^{(0)}$ via Eq. (2);
8:     Initialize cluster prompts $\mathbf{C}$ and assign to corresponding items via Eq. (3);
9:     Generate prompted item embeddings $\tilde{\mathbf{V}}^{(0)}$, applying normalization, self-gating and prompting function $\psi$ via Eq. (4)-(5);
10:     SBR model $f$ takes item set $\mathcal{S}$ and inital item embeddings $\tilde{\mathbf{V}}^{(0)}$, and return session embeddings $\mathbf{S}$ and updated item embeddings $\mathbf{V}$ via Eq. (6);
11:     for each item $v_{n}\in V$ do
12:        Compute recommendation scores $z_{k}$ and the probability $\hat{y}_{k}$ of item $v_{k}$ via Eq. (7)-(8);
13:     end for
14:     Compute loss $\mathcal{L}$ via Eq. (9) and update model parameters
15:  end for
@@ 3.3. Item Cluster-aware Prompt Learning
Table 2. Comparison of MRR@5 (M@5), Recall@5 (R@5), MRR@10 (M@10), and Recall@10 (R@10) on the three datasets. Our proposed models are highlighted in gray and the best results are highlighted in boldface.
Model Last.fm Xing Reddit
M@5 R@5 M@10 R@10 M@5 R@5 M@10 R@10 M@5 R@5 M@10 R@10
GRU4Rec 7.70 13.39 8.48 19.34 10.33 15.71 10.85 19.63 35.45 46.40 36.43 53.76
CLIP-GRU4Rec 10.07 17.24 10.94 23.87 11.12 16.90 11.69 21.19 40.21 50.47 41.06 56.85
NARM 5.50 9.74 6.15 14.65 9.66 15.41 10.27 20.02 33.08 44.48 34.09 52.04
CLIP-NARM 8.10 13.96 8.88 19.87 12.04 17.81 12.31 22.08 38.50 48.79 39.38 55.39
CORE 10.51 17.70 11.64 25.57 10.27 17.79 11.10 24.01 29.33 46.54 30.48 55.12
CLIP-CORE 10.67 18.11 11.58 24.99 12.65 19.59 13.35 24.86 31.33 47.56 32.37 52.05
SRGNN 9.02 14.66 9.76 20.33 14.03 19.37 14.49 22.59 38.04 47.75 38.88 54.05
CLIP-SRGNN 12.12 18.87 12.97 25.26 14.18 19.72 14.66 23.25 39.70 49.73 40.52 55.86
TAGNN 9.25 15.30 10.00 21.02 12.68 18.37 13.15 21.82 40.07 51.10 40.85 56.94
CLIP-TAGNN 12.82 19.77 13.70 26.50 14.70 21.07 15.25 25.18 41.22 52.01 42.02 57.96
GCSAN 4.47 8.69 5.03 12.99 13.02 16.82 13.71 24.07 36.03 44.13 36.81 50
CLIP-GCSAN 7.05 11.59 7.68 16.36 15.65 20.34 16.04 23.29 39.07 48.21 39.62 52.54
GCEGNN 10.81 17.32 11.64 23.54 12.83 19.41 13.45 24.02 38.17 48.38 39.07 55.08
CLIP-GCEGNN 11.05 17.60 11.89 23.94 14.36 21.58 14.96 26.00 37.93 48.19 38.84 55.08
SCL 9.99 14.90 10.49 20.31 10.58 15.88 11.23 22.84 31.25 40.43 32.66 46.54
CLIP-SCL 10.24 15.22 11.01 21.97 11.35 15.74 11.78 23.12 31.44 40.90 33.21 47.98
@@ 4. Experiments
  • RQ3: Does the Item Cluster-aware Prompt Learning module effectively and efficiently integrate item relationships into the SBR training?
@@ 4.1.2. Baselines
To evaluate the effectiveness of our approach, we compare its performance against a diverse set of well-established methods. The first five baselines focus soley on intra-session information, while the remaining four aim to enhance item relationship by incorporating additional contextual or gloabl information. We assess our method across various model architectures, including those based on RNNs, GNNs, attention mechanisms, and contrastive learning, to demonstrate its broad applicability and effectiveness.
@@
  • SCL (Shi et al., 2023) introduces a simplified contrastive learning framework that enhances item representations without requiring complex sample construction or data augmentation, improving recommendation performance efficiently. In this work, we implemented GCEGNN based SCL for evaluation.